# MADiff: Offline Multi-agent Learning with Diffusion Models

### v4

@@ 5.2 Compared Baselines and Metrics
For offline MARL experiments, we use the episodic return obtained in online rollout as the performance measure. We include MA-ICQ (Yang et al., 2021) and MA-CQL (Kumar et al., 2020) as baselines on all offline RL tasks. On MPE, we also include OMAR and MA-TD3+BC (Fujimoto and Gu, 2021) in baseline algorithms and use the results reported by Pan et al. (2022). On MA Mujoco, baseline results are adopted from Formanek et al. (2023). On SMAC, we include MADT (Meng et al., 2021b) as a sequence modeling baseline, while other baseline results are reported by Formanek et al. (2023). We implement independent behavior cloning (BC) as a naive supervised learning baseline.
@@ 5.3 Numerical Results
MATP on the NBA dataset. In Table 5.3, when comparing ADE and FDE, MADiff-C significantly outperforms the baseline; however, our algorithm only slightly beats baseline for $\text{minADE}_{20}$, and has higher $\text{minFDE}_{20}$. We suspect the reason is that Baller2Vec++ has a large prediction variance. When Baller2Vec++ only predicts one trajectory, a few players’ trajectories deviate from the truth so far that deteriorate the overall ADE and FDE. When allowing to sample 20 times and calculating the minimum ADE/FDE according to the ground truth, Baller2Vec++ can choose the best trajectory for every single agent, which makes $\text{minADE}_{20}$ and $\text{minFDE}_{20}$ significantly smaller than one-shot metrics. However, considering it may be not practical to select the best trajectories without access to the ground truth, MADiff-C is much more stable than Baller2Vec++. Predicted trajectories of MADiff-C and Baller2Vec++ are provided in the Appendix Section G.3.
@@ 5.4 Qualitative Analysis on Teammate modeling
In Figure 2 right, we demonstrate that such corrections of teammate modeling are common and can help agents make globally coherent behaviors. We sample 100 episodes with different initial states and define Consistent Ratio at some time step $t$ as the proportion of episodes in which the three agents make consistent planned trajectories. We plot the curve up to step $t=9$, which is approximately halfway through the episode length limit in MPE. The horizontal red line represents how many portions of the real rollout trajectories are consistent at step $t=9$. The interesting part is that the increasing curve reaches the red line before $t=9$, and ends up even higher. This indicates that the planned teammates’ trajectories are guiding the multi-agent interactions beforehand, which is a strong exemplar of the benefits of MADiff’s teammate modeling abilities. We also include visualizations of imagined teammate observation sequences in SMAC 3m task in the Appendix Section G.2.
@@ 5.5 Ablation Study
Our key argument is that the great coordination ability of MADiff is brought by the attention modules among individual agents’ diffusion networks. We validate this insight through a set of ablation experiments on MPE. We compare MADiff-D with independent DMs, i.e., each agent learns from corresponding offline data using independent U-Nets without attention. We denote this variant as MADiff-D-Ind. In addition, we also ablate the choice whether each agent should share parameters of their basic U-Net, noted as Share or NoShare. Without causing ambiguity, we omit the name of MADiff, and notate the different variants as D-Share, D-NoShare, Ind-Share, Ind-NoShare.
Figure 3: The average normalized score of MADiff ablation variants in MPE tasks. The mean and standard error are computed over 3 different seeds.
As is obviously observed in Figure 3, with attention modules, MADiff-D significantly exceeds that of the independent version on most tasks, justifying the importance of inter-agent attentions. The advantage of MADiff-D is more evident when the task becomes more challenging and the data becomes more confounded, e.g., results on World, where the gap between centralized and independent models is larger, indicating the difficulty of solving offline coordination with independently trained models. As for the parameter sharing choice, the performance of MADiff-D-Share and MADiff-D-NoShare is similar overall. Since MADiff-D-Share has fewer parameters, we prefer MADiff-D-Share, and use it as the default variant to be reported in Table 1. Another advantage of sharing U-Net parameters is that the trajectories of various agents can be batched together and fed through the network. This not only decreases sampling time but also renders it insensitive to an increasing number of agents. We provide a specific example in Appendix Section F.4.
@@ Appendix A Outline
In this appendix, we give the pseudocode of multi-agent planning and multi-agent trajectory prediction with MADiff model in Section B. In Section C, we demonstrate how multiple agents’ trajectories are modeled by MADiff during centralized control and decentralized execution in an example three-agent environment. In Section D, we give additional information on offline datasets, including how they are collected, violin plots of return distributions, and a minor issue of MPE dataset. In Section E, we briefly describe the implementation of baseline algorithms and links to related resources. In Section F, we provide details of the experiments, including the normalization used to compute the average score, the detailed network illustration unrolling each agent’s U-Net, crucial hyperparameters, and examples of wall-clock time and resources required for training and sampling from MADiff. In Section G, we demonstrate and analyze additional experimental results. We provide ablation results to support the effectiveness of teammate modeling in MADiff-D, show the quality of teammate modeling by MADiff-D on SMAC tasks, and visualize predicted multi-player trajectories by MADiff and the baseline algorithm on the NBA dataset.
Appendix B Algorithm
@@
Appendix C Illustration of Multi-agent Trajectory Modeling
@@
Appendix D Additional Information on offline datasets
D.1 MPE Datasets
For MPE experiments, we use datasets and a fork of environment***https://github.com/ling-pan/OMAR provided by OMAR [Pan et al., 2022]. They seem to be using an earlier version of MPE where agents can receive different rewards. For example, in the Spread task, team reward is defined using the distance of each landmark to its closest agent, which is the same for all agents. But when an agent collides with others, it will receive the team reward minus a penalty term. The collision reward has been brought into the team reward in the official repository since this commit***https://github.com/openai/multiagent-particle-envs/commit/6ed7cac026f0eb345d4c20232bafa1dc951c68e7. However, the fork provided by OMAR still uses a legacy version. For fair and proper comparisons, we use OMAR’s dataset and environment where all baseline models are trained and evaluated.
@@ D.3 SMAC Datasets
Appendix E Baseline Implementations
Here we briefly describe how the baseline algorithms are implemented. For MATP experiments, we use the implementation from the official repository of Baller2Vec++***https://github.com/airalcorn2/baller2vecplusplus. Baseline results on MPE datasets are borrowed from Pan et al. [2022]. According to their paper, they build all algorithms upon a modified version of MADDPG***https://github.com/shariqiqbal2810/maddpg-pytorch, which uses decentralized critics for all methods. Baselines on SMAC datasets are implemented by Formanek et al. [2023], and the performances are adopted from their reported benchmark results. The open-sourced implementation and hyperparameter settings can be found in the official repository***https://github.com/instadeepai/og-marl.
Appendix F Implementation Details
F.1 Score Normalization
@@
F.2 Detailed Network Architecture
@@ F.3 Hyperparameters
F.4 Computing Resources and Wall Time
@@
Appendix G Additional Experimental Results
G.1 Effectiveness of Teammate Modeling
@@
G.2 Teammate Modeling on SMAC Tasks
@@
G.3 Predicted Trajectory Visualization on NBA Dataset

### v5

@@ 5.2 Compared Baselines and Metrics
For offline MARL experiments, we use the episodic return obtained in online rollout as the performance measure. We include MA-ICQ (Yang et al., 2021) and MA-CQL (Kumar et al., 2020) as baselines on all offline RL tasks. On MPE, we also include OMAR and MA-TD3+BC (Fujimoto and Gu, 2021) in baseline algorithms and use the results reported by Pan et al. (2022). On MA Mujoco, baseline results are adopted from Formanek et al. (2023). On SMAC, we include MADT (Meng et al., 2021) as a sequence modeling baseline, while other baseline results are reported by Formanek et al. (2023). We implement independent behavior cloning (BC) as a naive supervised learning baseline.
@@ 5.3 Numerical Results
MATP on the NBA dataset. In Table 5.3, when comparing ADE and FDE, MADiff-C significantly outperforms the baseline; however, our algorithm only slightly beats baseline for $\text{minADE}_{20}$, and has higher $\text{minFDE}_{20}$. We suspect the reason is that Baller2Vec++ has a large prediction variance. When Baller2Vec++ only predicts one trajectory, a few players’ trajectories deviate from the truth so far that deteriorate the overall ADE and FDE. When allowing to sample 20 times and calculating the minimum ADE/FDE according to the ground truth, Baller2Vec++ can choose the best trajectory for every single agent, which makes $\text{minADE}_{20}$ and $\text{minFDE}_{20}$ significantly smaller than one-shot metrics. However, considering it may be not practical to select the best trajectories without access to the ground truth, MADiff-C is much more stable than Baller2Vec++. Predicted trajectories of MADiff-C and Baller2Vec++ are provided in the Appendix Section H.4.
@@ 5.4 Qualitative Analysis on Teammate modeling
In Figure 2 right, we demonstrate that such corrections of teammate modeling are common and can help agents make globally coherent behaviors. We sample 100 episodes with different initial states and define Consistent Ratio at some time step $t$ as the proportion of episodes in which the three agents make consistent planned trajectories. We plot the curve up to step $t=9$, which is approximately halfway through the episode length limit in MPE. The horizontal red line represents how many portions of the real rollout trajectories are consistent at step $t=9$. The interesting part is that the increasing curve reaches the red line before $t=9$, and ends up even higher. This indicates that the planned teammates’ trajectories are guiding the multi-agent interactions beforehand, which is a strong exemplar of the benefits of MADiff’s teammate modeling abilities. We also include visualizations of imagined teammate observation sequences in SMAC 3m task in the Appendix Section H.3.
@@ 5.5 Ablation Study
Our key argument is that the great coordination ability of MADiff is brought by the attention modules among individual agents’ diffusion networks. We validate this insight through a set of ablation experiments on MPE. We compare MADiff-D with independent DMs, i.e., each agent learns from corresponding offline data using independent U-Nets without attention. We denote this variant as MADiff-D-Ind. In addition, we also ablate the choice of whether each agent should share parameters of their basic U-Net, noted as Share or NoShare. Without causing ambiguity, we omit the name of MADiff, and notate the different variants as D-Share, D-NoShare, Ind-Share, Ind-NoShare.
Figure 3: The average normalized score of MADiff ablation variants in MPE tasks. The mean and standard error are computed over 5 different seeds.
As is obviously observed in Figure 3, with attention modules, MADiff-D significantly exceeds that of the independent version on most tasks, justifying the importance of inter-agent attentions. The advantage of MADiff-D is more evident when the task becomes more challenging and the data becomes more confounded, e.g., results on World, where the gap between centralized and independent models is larger, indicating the difficulty of solving offline coordination with independently trained models. As for the parameter sharing choice, the performance of MADiff-D-Share and MADiff-D-NoShare is similar overall. Since MADiff-D-Share has fewer parameters, we prefer MADiff-D-Share, and use it as the default variant to be reported in Table 1. Another advantage of sharing U-Net parameters is that the trajectories of various agents can be batched together and fed through the network. This not only decreases sampling time but also renders it insensitive to an increasing number of agents. We provide a specific example in Appendix Section G.4.
5.6 Limitations
Scalability to many agents. MADiff-D requires each agent to infer all teammates’ future trajectories, which is difficult and unnecessary in environments with a large number of agents. Although we have done experiments on a maximum number of 8 agents (SMAC 8m), MADiff-D is in general not suitable for scenarios with tens or hundreds of agents. A potential solution is to infer a latent representation of teammates’ trajectories.
Applicability in highly stochastic environments. Several theoretical and empirical studies (Paster et al., 2022; Brandfonbrener et al., 2022; Chen et al., 2021) have demonstrated that in offline RL, sequence modeling algorithms tend to underperform Q-learning-based algorithms in environments with high stochasticity. This is primarily because sequence modeling algorithms are more susceptible to high-reward offline trajectories that are achieved by chance. Since MADiff is a sequence modeling algorithm, it shares this weakness. To assess how much MADiff is affected by environmental stochasticity, we conducted experiments on the terran_5_vs_5 map in SMACv2 (Ellis et al., 2022). The design principle of SMACv2 is to add stochasticity to the original SMAC environment, including randomized initial positions and unit types. We conducted experiments under four settings: the original version, without position randomness, without unit type randomness, and without both kinds of randomness. MADiff performs worse than the Q-learning-based method only when both kinds of stochasticity are present. In all settings, MADiff outperforms the sequence modeling baseline. Detailed experimental settings and results can be found in Appendix Section H.1.
@@ 6 Conclusion
Acknowledgements
The SJTU team is partially supported by National Key R&D Program of China (2022ZD0114804), Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102) and National Natural Science Foundation of China (62322603, 62076161).
@@ Appendix A Outline
In this appendix, we provide a table to explain the main notations we used in Section B. In Section C, we give the pseudocode of multi-agent planning and multi-agent trajectory prediction with MADiff model. In Section D, we demonstrate how multiple agents’ trajectories are modeled by MADiff during centralized control and decentralized execution in an example three-agent environment. In Section E, we give additional information on offline datasets, including how they are collected, violin plots of return distributions, and a minor issue of MPE dataset. In Section F, we briefly describe the implementation of baseline algorithms and links to related resources. In Section G, we provide details of the experiments, including the normalization used to compute the average score, the detailed network illustration unrolling each agent’s U-Net, crucial hyperparameters, and examples of wall-clock time and resources required for training and sampling from MADiff. In Section H, we demonstrate and analyze additional experimental results. Specifically, we provide experiment results on SMACv2 to demonstrate how much MADiff is affected by environmental stochasticity. We also provide ablation results to support the effectiveness of teammate modeling in MADiff-D, show the quality of teammate modeling by MADiff-D on SMAC tasks, and visualize predicted multi-player trajectories by MADiff and the baseline algorithm on the NBA dataset.
Appendix B Notations
Table 3: List of main notations used in the paper.
Notation Description
$\mathcal{S},\mathcal{A},\Omega$ state, action, and local observation spaces
$\gamma$ the discounted factor
$N$ number of controlled agents
$s_{t}$ state at step $t$
$a^{i}_{t},o^{i}_{t}$ action and local observation of agent $i$ at environment step $t$
$\bm{a}_{t},\bm{o}_{t}$ joint action and observation of all agents at environment step $t$
$r(s,\bm{a})$ shared reward function
$\bm{\tau}$ joint trajectory of all agents
$\bm{y}(\bm{\tau})$ additional conditioning information
$\phi$ parameters of the inverse dynamics model
$\theta$ parameters of the diffusion model
$\tilde{x}^{i}_{k,t}$ noised observation of agent $i$ at diffusion step $k$ and environment step $t$
$\bm{\tilde{x}}_{k,t}$ noised joint observation at diffusion step $k$ and environment step $t$
$\hat{o}_{t}^{i}$ predicted observation of agent $i$ at environment step $t$
$\bm{\hat{\tau}}_{k}$ noised joint trajectory of all agents at diffusion step $k$
$h_{t}^{i}$ historical trajectory of agent $i$ up to environment step $t$
$\bm{h}_{t}$ historical joint trajectory of all agents up to environment step $t$
Appendix C Algorithm
@@
Appendix D Illustration of Multi-agent Trajectory Modeling
@@
Appendix E Additional Information on offline datasets
E.1 MPE Datasets
For MPE experiments, we use datasets and a fork of environment [^1] https://github.com/ling-pan/OMAR provided by OMAR [Pan et al., 2022]. They seem to be using an earlier version of MPE where agents can receive different rewards. For example, in the Spread task, team reward is defined using the distance of each landmark to its closest agent, which is the same for all agents. But when an agent collides with others, it will receive the team reward minus a penalty term. The collision reward has been brought into the team reward in the official repository since this commit [^2] https://github.com/openai/multiagent-particle-envs/commit/6ed7cac026f0eb345d4c20232bafa1dc951c68e7. However, the fork provided by OMAR still uses a legacy version. For fair and proper comparisons, we use OMAR’s dataset and environment where all baseline models are trained and evaluated.
@@ E.3 SMAC Datasets
Appendix F Baseline Implementations
Here we briefly describe how the baseline algorithms are implemented. For MATP experiments, we use the implementation from the official repository of Baller2Vec++ [^3] https://github.com/airalcorn2/baller2vecplusplus. Baseline results on MPE datasets are borrowed from Pan et al. [2022]. According to their paper, they build all algorithms upon a modified version of MADDPG [^4] https://github.com/shariqiqbal2810/maddpg-pytorch, which uses decentralized critics for all methods. Baselines on SMAC datasets are implemented by Formanek et al. [2023], and the performances are adopted from their reported benchmark results. The open-sourced implementation and hyperparameter settings can be found in the official repository [^5] https://github.com/instadeepai/og-marl.
Appendix G Implementation Details
G.1 Score Normalization
@@
G.2 Detailed Network Architecture
@@ G.3 Hyperparameters
G.4 Computing Resources and Wall Time
@@
Appendix H Additional Experimental Results
H.1 SMACv2 Experiments
To understand how much MADiff is affected by environmental stochasticity, we conducted experiments on the terran_5_vs_5 map in SMACv2 [Ellis et al., 2022]. SMACv2 is built upon SMAC with a focus on higher stochasticity. Specifically, in SMACv2, the unit types and agent start positions are randomized at the beginning of each episode. As each agent can only observe a nearby area, such randomness results in increased stochasticity in environment transitions. There are two different types of starting positions, reflect and surround. In reflect settings, the map is splitted into two sides. Allied units and enemy units are randomly and uniformly spawned on different sides. In surround settings, allied units are spawned at the center of the map, and enemy units are randomly stationed along the four diagonals. In terran_5_vs_5, there are three different unit types: marine, marauder, and medivac. The default sampling probabilities of these three types are 0.45, 0.45 and 0.1.
We design four settings with different degree of stochasicity: the original version, without position randomness (w/o PR), without unit type randomness (w/o TR), and without both kinds of randomness (w/o PR&TR). To reduce position randomness, we only use surrounding settings. Note that this does not mean the staring positions of all units are fixed, since enemy units are still randomized along the four diagonals. To remove unit type randomness, we set all units to be marines. The dataset for the original version is the terran_5_vs_5 Replay dataset from Formanek et al. [2023]. Datasets for other three stochasicity settings were collected by ourselves. We partially trained three MAPPO [^6] https://github.com/marlbenchmark/on-policy models in each setting. Each model was then used to collect 500 episodes, resulting in a dataset comprising 1500 episodes for each setting.
Three algorithms are benchmarked under these four settings: MAICQ, which represents the state-of-the-art in Q-learning-based algorithms; MADT, a representative multi-agent sequence modeling baseline; and MADiff-D. Results are presented in Table 8. We can see that MADiff-D performs worse than MAICQ only when both kinds of stochasticity are present. As the environmental randomness diminishes, MADiff-D’s performance gradually catches up with and surpasses MAICQ. In all settings, MADiff-D outperforms MADT.
Table 8: The average score on different settings of SMACv2 terran_5_vs_5. Shaded columns represent our method. The mean and standard error are computed over 3 different seeds.
Setting MAICQ MADT MADiff-D
Original $13.7\pm 1.7$ 8.2 $\pm$ 0.2 10.1 $\pm$ 0.8
w/o PR $16.0\pm 1.6$ 14.3 $\pm$ 0.8 $16.1\pm 0.3$
w/o TR $18.4\pm 0.5$ 14.6 $\pm$ 0.3 $18.6\pm 0.2$
w/o PR&TR 17.3 $\pm$ 0.3 16.8 $\pm$ 0.3 $18.5\pm 0.2$
H.2 Effectiveness of Teammate Modeling
@@
H.3 Teammate Modeling on SMAC Tasks
@@
H.4 Predicted Trajectory Visualization on NBA Dataset
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims made in the abstract and introduction accurately reflect our contributions.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We point out the limitations of our method in Section 5.6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: Our paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We illustrate our model architecture in Figure 1 and list important hyperparameters in Appendix Section G.3. We also provide the source code in supplementary materials.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide code, anonymous data download link, and necessary instructions in supplementary materials.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide the experimental details in Section 5.1, Section 5.2, and Appendix Section E, Section F, Section G.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We provide error bars in Table 1, Table 5.3, and Figure 3. Reported error bar is standard deviation calculated over trials with different random seeds.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide concrete instances of both training and sampling wall time and resources of our algorithm in Appendix Section G.4.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our paper conform with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: The proposed algorithm is a general solution for a wide range of offline multi-agent learning problems. In our opinion, there is no specific societal impact that should be stated explicitly.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cited and mentioned open-sourced implementation we used in Appendix Section F.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: Our paper does not publicly release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our paper does not involve human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our paper does not involve human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.